OpenStudy (anonymous):

PLEASE HELP ONLY 1 HOUR TO COMPLETE

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x – 14y = 6 –2x + 7y = –3

OpenStudy (anonymous):

go to mathway.com

OpenStudy (anonymous):

Thanks! The question is How many solutions are there to the following system of equations? 4x – 14y = 6 –2x + 7y = –3

OpenStudy (amoodarya):

multiply 2nd by 2 then add by 1st equation show me what happen

OpenStudy (nincompoop):

try solving it by multiplying the second equation by 2 then add them both

OpenStudy (nincompoop):

once you get the y value, then use it to solve for x using any of the two equations

OpenStudy (amoodarya):

4x – 14y = 6 –4x + 14y = –6 now add them

OpenStudy (anonymous):

Ok i gt it, i got infinitely many solutions

OpenStudy (amoodarya):

0=0 so has infinity solution you are in right track

OpenStudy (anonymous):

Solve the following system of equations. x + y = 3 2x + 2y = 6

OpenStudy (anonymous):

how about this one

OpenStudy (anonymous):

please help?